I am really struggling with this problem. I have come up with an average but am not sure where to go from here. The average I have is (2x+x+x-300)/3

Machine A can sort products twice as fast as Machine C and Machine B can sort products at a rate of 300 products fewer per minute than Machine C. If together the three machines have an average sorting rate of 2,150 products per minute, how many products per minute can Machine C sort?”

Answer by mananth(16946)

(4x-300)/3= 2150
4x-300=6450
4x=6450+300
4x=6750
x=1687.5
round off 1687